27).

The principal P is borrowed and the loans future value A at time t is given. Determine the loans simple interest rate r. PThe principal P is borrowed and the loans future value A at time t is given. Determine the loans simple interest rate r. P
Answer #1

Int = PTR

P = Principal Amount

T = Time period

R = Int Rate per anum

Int = FV - PV

Part A:

Int = FV - PV

= 6500 - 5000

= 1500

Int = PTR

1500 = 5000 * 4 * r

r = 1500 / 20000

= 0.075 i.e 7.5%

Part B:

Int = FV - PV

= 4380.63 - 4300

= 80.63

3 months = 0.25 Years

Int = PTR

80.63 = 4300 * 0.25 * r

r = 80.63 / 1075

= 0.075 i.e 7.5%

Pls comment, if any further assistance is required